# Asymmetries and Circumstellar Interaction in the Type II SN 2024bch

### v1

@@ 5.2 High-resolution spectra
As we show in Figure 9, the medium (GMOS) and high (PEPSI) resolution spectra around H taken on 2.9, 8, and 43 days after explosion give us critical information about the evolution of the ejecta and the physical characteristics of the surrounding CSM. On day 2.9 H can be fit with a Lorentzian profile with a Lorentzian full-width half-max (LFWHM) = 1700 km s-1 centered at km s-1 with electron scattering wings extending out to about 3000 km s-1. Additionally, there is a small bump just to the blue side of the narrow H that is emission from He II 6560 (see Smith et al., 2023). The narrow H is centered at 35 km s-1 and has a FWHM = 100 km s-1, with a possible P-Cygni absorption at km s-1 (although with the close He II it is difficult to tell). This narrow peak is presumably from radiatively heated CSM surrounding the SN that has yet to be hit by the SN shock, and the absorption minimum likely traces the progenitor wind speed. The resolution of the GMOS spectra is roughly 60 km s-1, so we are probably not fully resolving the line and overestimating the wind velocity. Of note is that typical RSG outflow speeds are only about 20 km s-1, so the speeds measured here are considerably faster.
@@
In the day 8 PEPSI spectrum we see two different components, neither exhibiting the Lorentzian shape as seen in early epochs. There is now an intermediate width feature that can be fit with a Gaussian having a FWHM = 800 km s-1 centered at km s-1 and showing a very prominent P-Cygni absorption component (as seen in the middle panel of Figure 9), and an additional narrow H emission feature with a FWHM = 35 km s-1 centered at approximately 10 km s-1. This narrow feature may have a very weak P-Cygni absorption with a minimum around km s-1, but it is at the noise level of the spectrum and hard to disentangle. Assuming it is a real feature, this would put the progenitor wind speed at 35 km s-1, much closer to RSG wind speeds, and similar to that found for SN 2024ggi (Shrestha et al., 2024a). Interestingly, this is much slower than the CSM speed seen a few days after explosion in SN 2023ixf (Smith et al., 2023). If the He II 6560 emission line is still present, it has been eroded by the strong H absorption. The broader P-Cygni absorption likely arises from gas that has been swept up by the shock, hence the speeds of up to 800–1000 km s-1, and cooled. This is also referred to as the cool dense shell (CDS). This absorption feature has a sharp minimum at km s-1 and wings extending out to roughly km s-1.
@@
In the bottom panel of Figure 9 we zoom in to compare the narrow components in the day 8 and 43 PEPSI spectra of SN 2024bch. By day 43 the faster, broad absorption line has disappeared, but the narrow emission line is now very pronounced with an obvious and deep P-Cygni absorption. The day 43 narrow line can be fit with a Gaussian with FWHM = 75 km s-1 centered at km s-1. The narrow emission component is coming from photoionized CSM, and the P-Cygni absorption, which has a minimum around km s-1 and extends out to roughly 100 km s-1, comes from the unshocked CSM along our line of sight. Similar features have been seen in interacting IIn supernovae observed in high resolution. For instance Smith & Andrews (2020) show that the IIn SN 2017hcc has narrow absorption persisting at least up to a year after explosion created from the non-shocked CSM. The later epochs of SN 2017hcc even reveal broader P-Cygni absorption as a result of the illumination of a wider range of absorption speeds along the line of sight, similar to what we see between day 8 and 43 in SN 2024bch. If we assume an average ejecta velocity of 7000 km s-1 as above, by 43 days the unshocked CSM would have to be at least 2.6 1015 cm away, and assuming a progenitor wind velocity of 35 km s-1 from the width of the narrow line, it would have to have been ejected about 25 years before explosion. This is an order of magnitude larger than the value estimated from the disappearance of the high ionization lines above. This discrepancy could be due to two distinct areas of CSM, one close in from the most recent mass loss of the progenitor creating the early narrow lines, and another extended region further out with possible asymmetries in the form of a disc or torus. The fact that we do see the narrow absorption features does suggest that at least some portion of the density enhancement if there is one, is along our line of sight.
@@
In the middle panel of Figure 9 we also compare PEPSI data of SN 2024bch on day 8 with that of SN 2023ixf on day 7.6 from Smith et al. (2023). While both SNe have similar red wing H emission, and the Gaussian FWHM for SN 2024bch is similar to the 900 km s-1 found for SN 2023ixf, the P-Cygni absorption is much broader and deeper in SN 2024bch, and spans a much larger range of velocities. This could indicate a different CSM geometry between the two objects. Smith et al. (2023) concluded that the CSM of SN 2023ixf was indeed asymmetric but was likely engulfed by the SN photosphere. While the possible asymmetric nature of the CSM for SN 2024bch will be discussed in more detail in section 6.3, the fact that we still see narrow P-Cygni H lines on day 43 does suggest that the CSM, at least along our line of sight, has not been completely swept up by the SN ejecta. If the CSM is asymmetric, it also requires that our line-of-sight has a vantage point looking through the dense CSM.
@@ 5.3 H Evolution
In Figure 10 we show the evolution of H in our lower resolution mostly Gemini/GMOS spectra from day 3 to day 144 normalized to the local continuum. While the early spectra show the narrow H discussed above, by day 22 a broad (7000 km s-1), blueshifted emission line starts to emerge. Blueshifted broad H lines are expected in the early spectra of Type II SNe as the redshifted receding side of the line is obscured by the opaque hydrogen envelope (Dessart & Hillier, 2005; Anderson et al., 2014). The blueshifted H peak migrates from km s-1 to km s-1 by day 56 where it stays until our last spectrum before Sun-occultation. Generally we expect the H line peak to move to zero velocity after the photospheric phase ends (Anderson et al., 2014), but this does not seem to be the case for SN 2024bch, at least by day 147. This could possibly be due to asymmetries in the supernova ejecta, an assumption further supported by the blueshifted peaks of the emerging [O I] doublet also at km s-1. The Type II SNe 2007it, 2017gmr, and more recently 2023ixf, all showed H and [O I] emission blueshifted by the same amount (Andrews et al., 2011, 2019; Ferrari et al., 2024). For the cases of SN 2017gmr and SN 2023ixf, spectropolarimetric observations further suggest intrinsic asymmetries in the SN ejecta (Nagao et al., 2019; Singh et al., 2024; Vasylyev et al., 2023; Shrestha et al., 2024b). This could very well be the case for SN 2024bch as we will explore more below.
Starting around 60–70 days, as the SN transitions from the plateau to the radioactive decay tail, an additional H emission feature begins to emerge around km s-1. Between day 80 and 125 this feature becomes very prominent and a bit flat topped, spanning between and km s-1, but by day 144 it has eroded and the center has moved redward to km s-1. In this last epoch, it is also possible that a similar feature is emerging on the red side around 4500 km s-1. Unfortunately, the SN becomes Sun constrained after this epoch, but future observations will be able to track this evolution. In the Type II SN PTF11iqb which also shows multi-peaked H and an emergence and eventually domination of a red side component, Smith et al. (2015) suggested that the asymmetric CSM was enveloped by the ejecta and was uncovered again as the photosphere receded. A similar situation could be occurring in SN 2024bch. Multi-component H like SN 2024bch exhibits has been observed in numerous other Type II SNe as well, and the postulated reason for these features is generally interaction with surrounding asymmetric CSM or an intrinsic asymmetry in the SN ejecta/explosion itself (Fransson et al., 2005; Andrews et al., 2017, 2019; Utrobin et al., 2021, for example). We will explore this more in depth in section 6.3.
@@ 5.4 NIR Spectra
Figure 11 shows the NIR spectral evolution from 2.7 to 60 days after explosion. The first spectrum is mainly featureless, except for narrow Pa and weak Br emission. Flash emission in the NIR is not commonly observed, whether due to lack of early spectra or lack of actual narrow lines. For example, in SN 2017gmr the NIR spectrum taken on day 2.8 was completely featureless (Andrews et al., 2019). In the following weeks, Pa broadens, and Pa emerges. By our last spectrum on day 60 strong Pa is seen, and all Paschen lines we observe have blueshifted peaks at km s-1. In Figure 12 we show Pa, Pa, and H in velocity space around day 60. The blueshifted peak is clear, and an additional component around km s-1 is seen emerging in Pa, a week or so before it becomes obvious in H as discussed above.
@@ 6.1 Model Comparison at Early Times
As seen in Figure 13, for SN 2024bch the persistence of the narrow Balmer, He II, and C IV lines until day 4 and likely beyond, and the overall shape of the continuum slope is most similar to the high mass-loss rate model r1w6 which uses an . At no time do we see N III/C III emission just blueward of He II 4685 that is seen in many other SNe with flash spectra (right panel of Figure 13 for example). It is very likely that the line was present but has faded by this epoch (see Tartaglia et al. (2024)). The very low resolution (R100) spectrum retrieved from TNS from day 1.3 (Balcon, 2024) does not have the resolution sufficient enough to differentiate individual lines in that region. Overall, the model comparison suggests that the mass-loss rate for SN 2024bch’s progenitor was likely between .
We have also compared our 2.9 day FLOYDS spectrum with models from Boian & Groh (2019) that simulate SN ejecta interaction with CSM roughly a day after explosion. They model three different progenitor types with different surface abundances: a solar abundance representing a low mass RSG, CNO-processed abundances from higher mass red-, blue-, and yellow- supergiants, and He-rich abundances corresponding to stars like LBVs or Wolf-Rayet stars. In the SN 2024bch spectrum we do not see any O VI, [O III], N III, or C III, but there is fairly strong H I, C IV, and He II 4685 emission. There are also weak but observable He II and N IV lines. As mentioned before, it is possible that some lines may have been present earlier and have faded by this epoch.
@@
Our earliest spectra were taken at 2.9 days after estimated explosion, almost two days after the models, so the comparison needs to be taken with some caution. Using the scaling relation as described in Boian & Groh (2019) though, we can calculate an L = 2.0 109 L⊙, and a range of = 3 – 7 10. This mass loss rate agrees fairly well with the comparison with the models from Dessart et al. (2017) discussed above, and is similar to estimated mass-loss rates from other Type II SNe as we will discuss in Section 6.2.
@@ 6.2 Comparison with other Flash SNe
Unfortunately we do not have spectra with suitable resolution before 2.9 days for SN 2024bch, but there are obvious similarities in He II 4685, C IV, and the Balmer lines between SN 2024bch and SN 2023ixf during this epoch. Similarly, there are no SN 2022jox spectra around day 3, but the day 1.5 spectrum also looks very similar to SN 2024bch. We also lack spectra for SN 2024bch between day 3.9 and 6.0 (shown in Figure 7) so it is unclear when the He II and C IV lines fully disappear, but SN 2023ixf and SN 2024bch seem quite similar around day 7 as well, with the notable exception being the obvious strong narrow P-Cygni absorption seen in H for SN 2024bch. This absorption feature is also seen in SN 2022jox 7.7 days after explosion. The flash features in SN 2024ggi evolve quickly, and are mostly gone (minus some Balmer line emission) by day 3 (Shrestha et al., 2024a), and overall do not match temporally with the other three SNe. Similarly, the -band lightcurves of SN 2022jox, SN 2023ixf, and SN 2024bch are all fairly similar, while SN 2024ggi has an extended plateau (Figure 4), and as shown in Figure 5 and mentioned in Shrestha et al. (2024a), the color of SN 2024ggi also reaches a maximum blue color much earlier than the other three.
@@ 6.3 Asymmetries in ejecta or CSM
Multiple features of the optical spectra of SN 2024bch lend themselves to the interpretation of asymmetries in either the CSM and/or the SN ejecta. In particular the blueshifted peaks of the Balmer lines along with those of [O I] in the nebular phase suggest that the explosion itself was asymmetric. “Horned” [O I] emission line profiles, where symmetric peaks are distributed around zero velocity, have been thoroughly studied in stripped envelope supernovae (SESN) (Modjaz et al., 2008; Taubenberger et al., 2009; Fang et al., 2022) and is attributed to aspherical ejecta geometries. These features seem to be less common in hydrogen rich SNe, but there are a few cases in the literature. Similar features were seen in SN 2023ixf (Ferrari et al., 2024) and SN 2017gmr (Andrews et al., 2019). These objects also had corresponding spectropolarimetry observations that confirm the asymmetric nature of the explosion (Nagao et al., 2019; Vasylyev et al., 2023; Singh et al., 2024; Shrestha et al., 2024b). In particular, Utrobin et al. (2021) and Fang et al. (2024) suggest that bipolar or toroidal oxygen ejecta can adequately reproduce the line profile of [O I] in SN 2023ixf.
@@
The multi-component H (and other hydrogen lines) seen in SN 2024bch may require additional asymmetries in the CSM. Asymmetric CSM interaction is not uncommon (Leonard et al., 2000; Smith et al., 2015; Andrews et al., 2010, 2017; Smith, 2017; Andrews & Smith, 2018; Andrews et al., 2019; Smith & Andrews, 2020; Smith et al., 2023), and the high binary fraction of massive stars (Sana et al., 2012) makes it likely that some objects may experience pre-SN binary interaction triggered by envelope inflation (Smith & Arnett, 2014). The appearance of the additional H feature around km s-1 that starts to emerge around day 60 and persists until our last epoch is likely being formed in the post-shock region between the forward and reverse shocks as the ejecta hits the CSM. That we only see a blue feature could easily be explained by a CSM geometry where from the observer’s viewing angle any back-side interaction is obscured by the SN ejecta. This hypothesis is further supported by the fact we still see narrow absorption from unshocked CSM in our nebular spectra. It is possible that an additional multi-component feature is emerging on the red side at a similar velocity in our last spectrum, but future observations are needed to confirm this. Of course one sided CSM, caused by various mechanisms could also be invoked, as it has been for previous objects (Smith et al., 2015; Thöne et al., 2017; Bilinski et al., 2018).

### v2

@@ 5.2 High-resolution spectra
As we show in Figure 9, the medium (GMOS) and high (PEPSI) resolution spectra around H taken on , and 43 days after explosion give us critical information about the evolution of the ejecta and the physical characteristics of the surrounding CSM. On day H can be fit with a Lorentzian profile with a Lorentzian full-width half-max (LFWHM) = 1700 km s-1 centered at km s-1 with electron scattering wings extending out to about 3000 km s-1. Additionally, there is a small bump just to the blue side of the narrow H that is emission from He II 6560 (see Smith et al., 2023). The narrow H is centered at 35 km s-1 and has a FWHM = 100 km s-1, with a possible P-Cygni absorption at km s-1 (although with the close He II it is difficult to tell). This narrow peak is presumably from radiatively heated CSM surrounding the SN that has yet to be hit by the SN shock, and the absorption minimum likely traces the progenitor wind speed. The resolution of the GMOS spectra is roughly 60 km s-1, so we are probably not fully resolving the line and overestimating the wind velocity. Of note is that typical RSG outflow speeds are only about 20 km s-1, so the speeds measured here are considerably faster.
@@
In the day PEPSI spectrum we see two different components, neither exhibiting the Lorentzian shape as seen in early epochs. There is now an intermediate width feature that can be fit with a Gaussian having a FWHM = 800 km s-1 centered at km s-1 and showing a very prominent P-Cygni absorption component (as seen in the middle panel of Figure 9), and an additional narrow H emission feature with a FWHM = 35 km s-1 centered at approximately 10 km s-1. This narrow feature may have a very weak P-Cygni absorption with a minimum around km s-1, but it is at the noise level of the spectrum and hard to disentangle. Assuming it is a real feature, this would put the progenitor wind speed at 35 km s-1, much closer to RSG wind speeds, and similar to that found for SN 2024ggi (Shrestha et al., 2024a) . Interestingly, this is much slower than the CSM speed measured a few days after explosion in SN 2023ixf (Smith et al., 2023). If the He II 6560 emission line is still present, it has been eroded by the strong H absorption. The broader P-Cygni absorption likely arises from gas that has been swept up by the shock, hence the speeds of up to 800–1000 km s-1, and cooled. This is also referred to as the cool dense shell (CDS). This absorption feature has a sharp minimum at km s-1 and wings extending out to roughly km s-1.
@@
In the bottom panel of Figure 9 we zoom in to compare the narrow components in the day and 43 PEPSI spectra of SN 2024bch. By day 43 the faster, broad absorption line has disappeared, but the narrow emission line is now very pronounced with an obvious and deep P-Cygni absorption. The day 43 narrow line can be fit with a Gaussian with FWHM = 75 km s-1 centered at km s-1. The narrow emission component is coming from photoionized CSM, and the P-Cygni absorption, which has a minimum around km s-1 and extends out to roughly 100 km s-1, comes from the unshocked CSM along our line of sight. Similar features have been seen in interacting IIn supernovae observed in high resolution. For instance Smith & Andrews (2020) show that the IIn SN 2017hcc has narrow absorption persisting at least up to a year after explosion created from the non-shocked CSM. The later epochs of SN 2017hcc even reveal broader P-Cygni absorption as a result of the illumination of a wider range of absorption speeds along the line of sight, similar to what we see between day 8 and 43 in SN 2024bch. If we assume an average ejecta velocity of 7000 km s-1 as above, by 43 days the unshocked CSM would have to be at least 2.6 1015 cm away, and assuming a progenitor wind velocity of 35 km s-1 from the width of the narrow line, it would have to have been ejected about 25 years before explosion. This is an order of magnitude larger than the value estimated from the disappearance of the high ionization lines above. This discrepancy could be due to two distinct areas of CSM, one close in from the most recent mass loss of the progenitor creating the early narrow lines, and another extended region further out with possible asymmetries in the form of a disc or torus. The fact that we do see the narrow absorption features does suggest that at least some portion of the density enhancement if there is one, is along our line of sight.
@@
In the middle panel of Figure 9 we also compare PEPSI data of SN 2024bch on day with the PEPSI data of SN 2023ixf on day 7.6 from Smith et al. (2023). While both SNe have similar red wing H emission, and the Gaussian FWHM for SN 2024bch is similar to the 900 km s-1 found for SN 2023ixf, the P-Cygni absorption is much broader and deeper in SN 2024bch, and spans a much larger range of velocities. This could indicate a different CSM geometry for the two objects. Smith et al. (2023) concluded that the CSM of SN 2023ixf was indeed asymmetric but was likely engulfed by the SN photosphere. While the possible asymmetric nature of the CSM for SN 2024bch will be discussed in more detail in section 6.3, the fact that we still see narrow P-Cygni H lines on day 43 does suggest that the CSM, at least along our line of sight, has not been completely swept up by the SN ejecta. If the CSM is asymmetric, it also requires that our line-of-sight has a vantage point looking through the dense CSM.
@@ 5.3 H Evolution
In Figure 10 we show the evolution of H in our lower resolution mostly Gemini/GMOS spectra from day 3 to day 144 normalized to the local continuum. While the early spectra show the narrow H discussed above, by day 22 a broad (7000 km s-1), blueshifted emission line starts to emerge. Blueshifted broad H lines are expected in the early spectra of Type II SNe as the redshifted receding side of the line is obscured by the opaque hydrogen envelope (Dessart & Hillier, 2005; Anderson et al., 2014a). The blueshifted H peak migrates from km s-1 to km s-1 by day 56 where it stays until our last spectrum before Sun-occultation. Generally we expect the H line peak to move to zero velocity after the photospheric phase ends (Anderson et al., 2014a), but this does not seem to be the case for SN 2024bch, at least by day 147. This could possibly be due to asymmetries in the supernova ejecta, an assumption further supported by the blueshifted peaks of the emerging [O I] doublet also at km s-1. The Type II SNe 2007it, 2017gmr, and more recently 2023ixf, all showed H and [O I] emission blueshifted by the same amount (Andrews et al., 2011, 2019; Ferrari et al., 2024). For the cases of SN 2017gmr and SN 2023ixf, spectropolarimetric observations further suggest intrinsic asymmetries in the SN ejecta (Nagao et al., 2019; Singh et al., 2024; Vasylyev et al., 2023; Shrestha et al., 2024b). This could very well be the case for SN 2024bch as we will explore more below.
Starting around 60–70 days, as the SN transitions from the plateau to the radioactive decay tail, an additional H emission feature begins to emerge around km s-1. Between day 80 and 125 this feature becomes very prominent and a bit flat topped, spanning between and km s-1, but by day 144 it has eroded and the center has moved redward to km s-1. In this last epoch, it is also possible that a similar feature is emerging on the red side around 4500 km s-1. Unfortunately, the SN becomes Sun constrained after this epoch, but future observations will be able to track this evolution. In the Type II SN PTF11iqb which also shows multi-peaked H and an emergence and eventually domination of a red side component, Smith et al. (2015) suggested that the asymmetric CSM was enveloped by the ejecta and was uncovered again as the photosphere receded. A similar situation could be occurring in SN 2024bch. Multi-component H like SN 2024bch exhibits has been observed in numerous other Type II SNe as well, and the postulated reason for these features is generally interaction with surrounding asymmetric CSM or an intrinsic asymmetry in the SN ejecta/explosion itself (Fransson et al., 2005; Andrews et al., 2017, 2019; Utrobin et al., 2021; Pessi et al., 2023, for example). We will explore this more in depth in section 6.3.
@@ 5.4 NIR Spectra
Figure 11 shows the NIR spectral evolution from to 60 days after explosion. The first spectrum is mainly featureless, except for narrow Pa and weak Br emission. Flash emission in the NIR is not commonly observed, whether due to lack of early spectra or lack of actual narrow lines. For example, in SN 2017gmr the NIR spectrum taken on day 2.8 was completely featureless (Andrews et al., 2019). In the following weeks, Pa broadens, and Pa emerges. By our last spectrum on day 60 strong Pa is seen, and all Paschen lines we observe have blueshifted peaks at km s-1. In Figure 12 we show Pa, Pa, and H in velocity space around day 60. The blueshifted peak is clear, and an additional component around km s-1 is seen emerging in Pa, a week or so before it becomes obvious in H as discussed above.
@@ 6.1 Model Comparison at Early Times
As seen in Figure 13, for SN 2024bch the persistence of the narrow Balmer, He II, and C IV lines until day 3.8 and likely beyond, and the overall shape of the continuum slope is most similar to the high mass-loss rate model r1w6 which uses an . At no time do we see N III/C III emission just blueward of He II 4685 that is seen in many other SNe with flash spectra (right panel of Figure 13 for example). It is very likely that the line was present but has faded by this epoch (see Tartaglia et al. (2024)). The very low resolution (R100) spectrum retrieved from TNS from day (Balcon, 2024) does not have the resolution sufficient enough to differentiate individual lines in that region. Overall, the model comparison suggests that the mass-loss rate for SN 2024bch’s progenitor was likely between .
We have also compared our day FLOYDS spectrum with models from Boian & Groh (2019) that simulate SN ejecta interaction with CSM roughly a day after explosion. They model three different progenitor types with different surface abundances: a solar abundance representing a low mass RSG, CNO-processed abundances from higher mass red-, blue-, and yellow- supergiants, and He-rich abundances corresponding to stars like LBVs or Wolf-Rayet stars. In the SN 2024bch spectrum we do not see any O VI, [O III], N III, or C III, but there is fairly strong H I, C IV, and He II 4685 emission. There are also weak but observable He II and N IV lines. As mentioned before, it is possible that some lines may have been present earlier and have faded by this epoch.
@@
Our earliest spectra were taken at days after estimated explosion, almost two days after the models, so the comparison needs to be taken with some caution. Using the scaling relation as described in Boian & Groh (2019) though, we can calculate an L = 2.0 109 L⊙, and a range of = 3 – 7 10. This mass loss rate agrees fairly well with the comparison with the models from Dessart et al. (2017) discussed above, and is similar to estimated mass-loss rates from other Type II SNe as we will discuss in Section 6.2.
@@ 6.2 Comparison with other Flash SNe
Unfortunately we do not have spectra with suitable resolution before days for SN 2024bch, but there are obvious similarities in He II 4685, C IV, and the Balmer lines between SN 2024bch and SN 2023ixf during this epoch. Similarly, there are no SN 2022jox spectra around day 3, but the day 1.5 spectrum also looks very similar to SN 2024bch. We also lack spectra for SN 2024bch between day (shown in Figure 7) so it is unclear when the He II and C IV lines fully disappear, but SN 2023ixf and SN 2024bch seem quite similar around day 7 as well, with the notable exception being the obvious strong narrow P-Cygni absorption seen in H for SN 2024bch. This absorption feature is also seen in SN 2022jox 7.7 days after explosion. The flash features in SN 2024ggi evolve quickly, and are mostly gone (minus some Balmer line emission) by day 3 (Shrestha et al., 2024a), and overall do not match temporally with the other three SNe. Similarly, the -band lightcurves of SN 2022jox, SN 2023ixf, and SN 2024bch are all fairly similar, while SN 2024ggi has an extended plateau (Figure 4), and as shown in Figure 5 and mentioned in Shrestha et al. (2024a), the color of SN 2024ggi also reaches a maximum blue color much earlier than the other three.
@@ 6.3 Asymmetries in ejecta or CSM
Multiple features of the optical spectra of SN 2024bch lend themselves to the interpretation of asymmetries in either the CSM and/or the SN ejecta. In particular the blueshifted peaks of the Balmer lines along with those of [O I] in the nebular phase suggest that the explosion itself was asymmetric. “Horned” [O I] emission line profiles, where symmetric peaks are distributed around zero velocity, have been thoroughly studied in stripped envelope supernovae (SESN) (Modjaz et al., 2008; Taubenberger et al., 2009; Fang et al., 2022, 2024a) and is attributed to aspherical ejecta geometries. These features seem to be less common in hydrogen rich SNe, but there are a few cases in the literature. Similar features were seen in SN 2023ixf (Ferrari et al., 2024) and SN 2017gmr (Andrews et al., 2019). These objects also had corresponding spectropolarimetry observations that confirm the asymmetric nature of the explosion (Nagao et al., 2019; Vasylyev et al., 2023; Singh et al., 2024; Shrestha et al., 2024b). In particular, Utrobin et al. (2021) and Fang et al. (2024b) suggest that bipolar or toroidal oxygen ejecta can adequately reproduce the line profile of [O I] in SN 2023ixf.
@@
The multi-component H (and other hydrogen lines) seen in SN 2024bch may require additional asymmetries in the CSM. Asymmetric CSM interaction is not uncommon (Leonard et al., 2000; Smith et al., 2015; Andrews et al., 2010, 2017; Smith, 2017; Andrews & Smith, 2018; Andrews et al., 2019; Smith & Andrews, 2020; Smith et al., 2023), and the high binary fraction of massive stars (Sana et al., 2012) makes it likely that some objects may experience pre-SN binary interaction triggered by envelope inflation (Smith & Arnett, 2014; Soker, 2013). The appearance of the additional H feature around km s-1 that starts to emerge around day 60 and persists until our last epoch is likely being formed in the post-shock region between the forward and reverse shocks as the ejecta hits the CSM. That we only see a blue feature could easily be explained by a CSM geometry where from the observer’s viewing angle any back-side interaction is obscured by the SN ejecta. This hypothesis is further supported by the fact we still see narrow absorption from unshocked CSM in our nebular spectra. It is possible that an additional multi-component feature is emerging on the red side at a similar velocity in our last spectrum, but future observations are needed to confirm this. Of course one sided CSM, caused by various mechanisms could also be invoked, as it has been for previous objects (Smith et al., 2015; Thöne et al., 2017; Bilinski et al., 2018).